OpenStudy (anonymous):

Write a polynomial function of least degree with integral coefficient that has the given zeros. 6, 2+2i

hero (hero):

Basically, 6 and 2 + 2i are two solutions of the polynomial function when f(x) = 0. In other words, x = 6 x = 2 + 2i But remember that solutions involving complex numbers occur as conjugate pairs so x = 2 - 2i is also a solution. Now next, we can move all the numbers to one side so that each equation is equal to zero: x - 6 = 0 x - 2 - 2i = 0 x - 2 + 2i = 0 And next, we can reference the zero product property to show that (x - 6)(x - 2 - 2i)(x - 2 + 2i) = 0 From here, we can multiply the left hand side and we will have the polynomial of simplest form that includes all the given zeroes.

hero (hero):

So if you multiply: (x - 6)(x - 2 - 2i)(x - 2 + 2i) you will get the polynomial function of least degree.

hero (hero):

@DSoul_64 does any of what I have shown so far make any sense to you at all?

OpenStudy (anonymous):

A little

hero (hero):

Where exactly are you confused? Would you mind explaining what you understand, or more importantly what you do not understand?

OpenStudy (anonymous):

Mostly when to multiply the (x-2+2i) and (x-2-2i)

OpenStudy (anonymous):

Be back in 30 min

hero (hero):

Suppose we were working with real numbers only. The question states that the zeroes of a polynomial are 2, 3, and 4. You know that this means the solutions of the polynomial where f(x) = 0 occur where x = 2 x = 3 x = 4 You also know that if you move everything to one side you will have x - 2 = 0 x - 3 = 0 x - 4 = 0 Now, to get the polynomial that includes these x values, you have to write out this equation in reference to the zero product property: (x - 2)(x - 3)(x - 4) = 0 Then multiply the binomials on the left side to get: x^3 - 9x^2 + 26x - 24 = f(x)